# Two Tales of Persona in LLMs: A Survey of Role-Playing and Personalization

### v2

@@ 1 Introduction
for tree= forked edges, draw=edgeColor, thick, font=, fill=tnodeColor, rectangle, rounded corners=4pt, text=textColor, blur shadow=shadow scale=0.95, shadow xshift=.5ex, shadow yshift=-.5ex, shadow opacity=0.25, edge=-, draw=edgeColor, line width=2pt, grow=0, child anchor=west, parent anchor=east, anchor=west, align=center, l sep+=0.3cm, s sep+=0.1cm, , root/.style= fill=Gray!15, font=, rounded corners=6pt, text width=4cm, /tikz/align=center, inner sep=6pt, tnode2_1/.style=fill=Gray!15, font=, text width=12cm, /tikz/align=center, inner sep=6pt, tnode2_2/.style=fill=Gray!15, font=, text width=11cm, /tikz/align=center, inner sep=6pt, tnode2_3/.style=fill=Gray!15, font=, text width=12cm, /tikz/align=center, inner sep=6pt, tnode3_1/.style=fill=Gray!15, font=, text width=12cm, /tikz/align=center, inner sep=6pt, tnode3_2/.style=fill=Gray!15, font=, text width=11cm, /tikz/align=center, inner sep=6pt, tnode3_3/.style=fill=Gray!15, font=, text width=12cm, /tikz/align=center, inner sep=6pt, [Taxonomy, root [LLM Personality Evaluation Section 4, tnode2_1 [BigFive; MBTI; etc, tnode2_2] ] [LLM Personalization Section 3, tnode2_1 [Dialogue Section 3.5, tnode2_2 [User Persona Modeling, tnode2_3] [Task-Oriented Modeling, tnode2_3] ] [Healthcare Section 3.4, tnode2_2] [Education Section 3.3, tnode2_2] [Search Section 3.2, tnode2_2] [Recommendation Section 3.1, tnode2_2] ] [LLM Role-Playing Section 2, tnode3_1 [Emergent Behaviors Section 2.3, tnode3_2 [Destructive Behavior, tnode3_3] [Conformity Behavior, tnode3_3] [Voluntary Behavior, tnode3_3] ] [Role-Playing Schema Section 2.2, tnode3_2 [Multi-Agent, tnode3_3] [Single-Agent, tnode3_3] ] [Environments Section 2.1, tnode3_2 [LLM-as-Evaluator Section 2.1.4, tnode3_3] [Medical Application Section 2.1.3, tnode3_3] [Game Section 2.1.2, tnode3_3] [Software Development Section 2.1.1, tnode3_3] ] ] ]
@@
Figure 2: The taxonomy of LLM role-playing and LLM personalization.
@@
In this paper, we aim to close the gap by offering a comprehensive survey and a systematic categorization of existing studies. Specifically, we divide current research into two main streams, namely LLM Role-Playing and LLM Personalization, as illustrated in Figure 1. The primary distinction is that in role-playing, the persona belongs to the LLM, while in personalization, the persona belongs to the user. The definitions are detailed below.
@@
Our taxonomy is illustrated in Figure 2. We first introduce LLM role-playing Section 2, followed by LLM personalization Section 3. Next, we provide an overview of evaluation methods Section 4 assessing whether the personality of LLMs (e.g., personality traits or psychological behaviors) accurately aligns with expected personas after the adaptation (i.e., for role-playing LLMs that act according to assigned personas and personalized LLMs that fit user personas). Lastly, we highlight current challenges and future directions Section 5. A comprehensive list of benchmarks and datasets is provided in the Appendix.
@@ 2 LLM Role-Playing
Figure 3: An illustration of five LLM role-playing environments: Software Development Section 2.1.1, Game Section 2.1.2, Medical Application Section 2.1.3, and LLM as Evaluators Section 2.1.4. For each environment, we provide a simple scenario with a task description (red-bordered) and relevant personas (i.e., roles; blue-bordered). The dashed rectangle represents an example LLM role-playing prompt template. In addition to the above environments, past research also proposes general frameworks applicable to different environments Section 5.1.
@@ 3.1 Personalized Recommendation
Existing studies explore various prompting methods for using LLMs in recommendation systems. Li et al. (2023a) develop a method for efficient incorporation of users’ personal information. Li et al. (2023b) combine aspect extraction with aspect-based recommendations via LLMs prompt tuning. Chen et al. (2022) generate personalized chit-chat to enhance recommendation. Focusing on the framework design, Yang et al. (2023b) present a novel LLM fine-tuning recommendation system. Chu et al. (2023) merge different recommendation systems to effectively integrating the commonsense and reasoning abilities of LLMs. Hu et al. (2024) propose a sequential recommendation framework to preserve fine-grained item textual information.
@@ 4 LLM Personality Evaluation
5 Limitations and Future Directions
@@ 5.3 Lack of Datasets and Benchmarks
For LLM role-playing, several tasks lack suitable datasets with specific formats Ahn et al. (2024) and environmental information (e.g., game environments require information about configurations and tools). For personalized dialogue generation, user persona modeling lacks contradictory persona datasets and multimodal persona datasets that would more accurately represent real human behaviors (Kim et al., 2024b; Ahn et al., 2023). Furthermore, LLM personalization faces a scarcity of high-quality personal data for model development due to privacy concerns, hindering a thorough evaluation of different personalization methods. In addition, existing benchmarks for both LLM role-playing and personalization are relatively limited, lacking comprehensive evaluations across various dimensions (Chang et al., 2023). Therefore, expanding datasets and benchmarks for specialized environments and personal information under privacy protection is an important future direction.
@@ 5.4 Bias
While a large number of studies focus on enhancing end-task performance, fewer works explore the biases induced by role-playing and personlization in LLMs. In this context, Gupta et al. (2023), as one of the first studies, highlights the deep-seated stereotypical biases found in LLMs assigned with socio-demographic personas. For personalized LLM recommendation systems, biases can be observed due to item popularity or item positions in the prompts Hou et al. (2024). Empirically, Dorner et al. (2023) also reveal the presence of agree bias in LLMs – a tendency to agree with both true and false content, regardless of the actual facts. In sum, there exists ample room for investigating and mitigating different classes of biases in the context of LLM role-playing and personalization.
@@ 5.5 Safety and Privacy
Past research has shown safety issues in LLM role-playing and personalization. Jin et al. (2024a) and Shah et al. (2023) successfully manipulate LLMs to perform jailbreak collaboratively. Deshpande et al. (2023) also show that assigning personas to LLMs aid in jailbreaking. Negative behaviors in LLM role-playing are also demonstrated by Chen et al. (2023c) and Li et al. (2024a). Further, Deshpande et al. (2023) find that LLMs consistently exhibit toxicity in a range of topics when assigned personas. These works demonstrate the discovery of unsafe problems, indicating an urgent need and more efforts to prevent potential exploits.
@@ Acknowledgements
We thank Yu-Ching Hsu and Jia-Yin Foo from National Taiwan University for their assistance and discussion. This work is financially supported by the National Science and Technology Council (NSTC), Taiwan, under Grants 112-2223-E-002-012-MY5 and 111-$2$-E-002-013-MY3, and from the Featured Area Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education (113L900901/113L900902/113L900903).
@@ Appendix A Web
Liu et al. (2023) Beauty Amazon Prompting Rating Prediction, Sequential Recommendation, Direct Recommendation, Explanation Generation and Review Summarization
@@
Chen et al. (2022) News MIND Wu et al. (2020), Reddit Fine-tuning with weak labels Direct Recommendation
Table 2: An overview of existing research in recommendation. Following the classification of Liu et al. (2023), we classify recommendation systems into five types: rating prediction, sequential recommendation, explanation Generation, and review generation, and direct recommendation.
@@
User Persona PersonaChat Zhang et al. (2018a) 11,907 164,356 1
@@
Table 4: A list of commonly used datasets for ToD modeling and user persona modeling. Among them, different versions of MultiWOZ Budzianowski et al. (2018); Ramadan et al. (2018); Eric et al. (2020); Zang et al. (2020) and PersonaChat Zhang et al. (2018a) are the most commonly used. Updated versions of MultiWOZ improve in several aspects: data quality, dialogue complexity, schema and ontology updates, and dataset sizes. PersonaChat contains various persona profiles, consisting of background, preferences, and personality traits. These profiles enable the modeling of coherent and contextual multi-turn diverse dialogue scenarios. For applications in user persona modeling, Tu et al. (2023) match individuals with persona-compatible virtual supporters and introduces the MBTI-S2Conv dataset, containing conversations between characters with distinct profiles. Lotfi et al. (2024) and Han et al. (2024) both propose synthetic datasets related to the Big Five personality.
@@
Table 6: An example of user persona modeling Section 3.5 from Persona-Chat dataset Zhang et al. (2018b).

### v3

@@ 1 Introduction
for tree= forked edges, draw=edgeColor, thick, font=, fill=tnodeColor, rectangle, rounded corners=4pt, text=textColor, blur shadow=shadow scale=0.95, shadow xshift=.5ex, shadow yshift=-.5ex, shadow opacity=0.25, edge=-, draw=edgeColor, line width=2pt, grow=0, child anchor=west, parent anchor=east, anchor=west, align=center, l sep+=0.3cm, s sep+=0.3cm, , root/.style= fill=Gray!15, font=, rounded corners=6pt, text width=4cm, /tikz/align=center, inner sep=6pt, tnode2/.style=fill=Gray!15, font=, text width=11cm, /tikz/align=center, inner sep=6pt, tnode3/.style=fill=Gray!15, font=, text width=9cm, /tikz/align=center, inner sep=6pt, tnode4/.style=fill=Gray!15, font=, text width=10cm, /tikz/align=center, inner sep=6pt, cnode4/.style=fill=Yellow!15, font=, edge=dotted, text width=22.7cm, align=left, cnode5/.style=fill=Yellow!15, font=, edge=dotted, text width=11.5cm, align=left, [Taxonomy, root [LLM Personality Evaluation Section 4, tnode2 [Big Five; MBTI; etc, tnode3 [ Jiang et al. (2023); Sorokovikova et al. (2024);
 Pan and Zeng (2023); Song et al. (2024);  PsychoBench (Huang et al., 2023), cnode4] ] ] [LLM Personalization Section 3, tnode2 [Dialogue Section 3.5, tnode3 [User Persona Modeling, tnode4 [ $\mathcal{P}^{2}$ (Liu et al., 2020);  Kim et al. (2024b);
 CLV (Tang et al., 2023b)
 – Dataset –
 PersonaChat (Zhang et al., 2018a);
 ConvAI2 PersonaChat (Dinan et al., 2019), cnode5] ] [ToD Modeling, tnode4 [ RefGPT (Yang et al., 2023a);
 ProToD (Hu et al., 2023);
 DSP (Li et al., 2024c)
 – Dataset –
 MutiWoz (Budzianowski et al., 2018);
 SGD (Rastogi et al., 2020), cnode5] ] ] [Healthcare Section 3.4, tnode3 [ openCHA (Abbasian et al., 2024a);  CHA (Abbasian et al., 2024c);
 MaLP (Zhang et al., 2024a);  HealthLLM (Jin et al., 2024b)
 – Dataset –
 MedDialog (Zeng et al., 2020);  IMCS-21 (Chen et al., 2022b);  iCliniq (Li et al., 2023d), cnode4], ] [Education Section 3.3, tnode3 [ HumSum Shehata et al. (2023);  EduChat (Dan et al., 2023);  Park et al. (2024);
 Kasneci et al. (2023);  Wang et al. (2024b), cnode4] ] [Search Section 3.2, tnode3 [ CoPS (Zhou et al., 2024);  Spatharioti et al. (2023);  Ziems et al. (2023);
 Sharma et al. (2024);  Baek et al. (2024), cnode4] ] [Recommendation Section 3.1, tnode3 [ PALR (Yang et al., 2023b);  Dai et al. (2023);  ONCE (Liu et al., 2023b)
 Christakopoulou et al. (2023);  BookGPT (Zhiyuli et al., 2023);
 – Dataset –
 MovieLens (Harper and Konstan, 2015), cnode4] ] ] [LLM Role-Playing Section 2, tnode2 [Emergent Behaviors Section 2.3, tnode3 [Destructive Behavior, tnode4 [ CAMEL (Li et al., 2024a);
 Deshpande et al. (2023);
 Gupta et al. (2023), cnode5] ] [Conformity Behavior, tnode4 [ MedAgents (Tang et al., 2023a);
 Fu et al. (2023), cnode5] ] [Voluntary Behavior, tnode4 [ AgentVerse (Chen et al., 2023c);
 MetaGPT (Hong et al., 2023a), cnode5] ] ] [Role-Playing Schema Section 2.2, tnode3 [Multi-Agent, tnode4 [ ChatDev (Qian et al., 2023);
 MedAgents (Tang et al., 2023a), cnode5] ] [Single-Agent, tnode4 [ Voyager (Wang et al., 2023a);
 MindAct (Deng et al., 2024), cnode5] ] ] [Environments Section 2.1, tnode3 [LLM-as-Evaluator Section 2.1.4, tnode4 [ DRPE (Wu et al., 2023b);
 ChatEval (Chan et al., 2023), cnode5] ] [Medical Application Section 2.1.3, tnode4 [ Wu et al. (2023a);
 MedAgents (Tang et al., 2023a), cnode5] ] [Game Section 2.1.2, tnode4 [ Generative Agents (Park et al., 2023);
 Humanoid Agent (Wang et al., 2023d);
 Fu et al. (2023);
 Voyager (Wang et al., 2023a), cnode5] ] [Software Development Section 2.1.1, tnode4 [ Dong et al. (2023);
 ChatDev (Qian et al., 2023);
 MetaGPT (Hong et al., 2023a)
 – Dataset –
 SoftwareDev (Hong et al., 2023a);
 SRDD (Qian et al., 2023), cnode5] ] ] ] ]
@@
Figure 2: The taxonomy of LLM role-playing and LLM personalization (representative works shown only).
@@
In this paper, we aim to close the gap by offering a comprehensive survey and a systematic categorization of existing studies. Specifically, we divide current research into two main streams, namely LLM Role-Playing and LLM Personalization, as illustrated in Figure 1. The primary distinction is that in role-playing, the persona belongs to the LLM, while in personalization, the persona belongs to the user. Further, the literature on role-playing mainly focuses on the tasks (i.e., how LLMs with role-playing can achieve better performance). In contrast, the literature of personalization primarily focuses on the users (i.e., how to satisfy users’ expectations and meet their needs). It is noteworthy that both of role-playing and personalization can be goals in the same scenario, but serve different purposes and are driven by different aspects. The definitions are detailed below.
@@
Our taxonomy is illustrated in Figure 2. We first introduce LLM role-playing Section 2, followed by LLM personalization Section 3. Next, we provide an overview of evaluation methods Section 4 assessing whether the personality of LLMs (e.g., personality traits or psychological behaviors) accurately aligns with expected personas after the adaptation (i.e., for role-playing LLMs that act according to assigned personas and personalized LLMs that fit user personas). Lastly, we highlight current challenges and future directions Section 5. We hope that this taxonomy could serve as a useful guideline for researchers to easily target the tasks/scenarios of interests, and swiftly pinpoint their current position in the field.
Figure 3: An illustration of four LLM role-playing environments: Software Development Section 2.1.1, Game Section 2.1.2, Medical Application Section 2.1.3, and LLM as Evaluators Section 2.1.4. For each environment, we provide a simple scenario with a task description (red-bordered) and relevant personas (i.e., roles; blue-bordered). The dashed rectangle represents an example LLM role-playing prompt template. In addition to the above environments, past research also proposes general frameworks applicable to different environments Section 5.1.
@@ 3.1 Personalized Recommendation
Existing studies explore various prompting methods for using LLMs in recommendation systems. Li et al. (2023a) develop a method for efficient incorporation of users’ personal information. Li et al. (2023b) combine aspect extraction with aspect-based recommendations via LLMs prompt tuning. Chen et al. (2022a) generate personalized chit-chat to enhance recommendation. Focusing on the framework design, Yang et al. (2023b) present a novel LLM fine-tuning recommendation system. Chu et al. (2023) merge different recommendation systems to effectively integrating the commonsense and reasoning abilities of LLMs. Hu et al. (2024) propose a sequential recommendation framework to preserve fine-grained item textual information.
@@ 4 LLM Personality Evaluation
5 Challenges and Future Directions
@@ 5.3 Lack of Datasets and Benchmarks
For LLM role-playing, several tasks lack suitable datasets with specific formats Ahn et al. (2024) and environmental information (e.g., game environments require information about configurations and tools). For personalized dialogue generation, user persona modeling lacks contradictory persona datasets and multimodal persona datasets that would more accurately represent real human behaviors (Kim et al., 2024b; Ahn et al., 2023). Furthermore, LLM personalization faces a scarcity of high-quality personal data for model development due to privacy concerns, hindering a thorough evaluation of personalization methods. In addition, existing benchmarks for both LLM role-playing and personalization are relatively limited, lacking comprehensive evaluations across various dimensions (Chang et al., 2023; Samuel et al., 2024). Therefore, expanding datasets and benchmarks for specialized environments and personal information under privacy protection is an important next step.
@@ 5.4 Bias
While a large number of studies focus on enhancing end-task performance, fewer works explore the biases induced by role-playing and personalization in LLMs. In this context, Gupta et al. (2023), as one of the first studies, highlight the deep-seated stereotypical biases found in LLMs assigned with socio-demographic personas. Additionally, Zhao et al. (2024) find that applying role-play often increases the overall likelihood of generating stereotypical and harmful outputs. For personalized LLM recommendation systems, biases can be observed due to item popularity or item positions in the prompts Hou et al. (2024). Empirically, Dorner et al. (2023) also reveal the presence of agree bias in LLMs – a tendency to agree with both true and false content, regardless of the actual facts. In sum, there exists ample room for investigating and mitigating different classes of biases in the context of LLM role-playing and personalization.
@@ 5.5 Safety and Privacy
Past research has shown safety issues in LLM role-playing and personalization. Jin et al. (2024a) and Shah et al. (2023) successfully manipulate LLMs to perform jailbreak collaboratively. Deshpande et al. (2023) also show that assigning personas to LLMs aid in jailbreaking. Negative behaviors in LLM role-playing are also demonstrated by Chen et al. (2023c) and Li et al. (2024a). Further, Deshpande et al. (2023) find that LLMs consistently exhibit toxicity in a range of topics when assigned personas, and  Vijjini et al. (2024) show that LLMs suffer from personalization bias when they are personalized for the user’s demographic. These works demonstrate the discovery of unsafe problems, indicating an urgent need and more efforts to prevent potential exploits.
@@ 7 Conclusion
Limitations
For the evaluation metric, as the literature, even within the same scope, addresses various sub-tasks and employs different corresponding evaluation metrics, or proposes their own ones (e.g., persona accuracy, task success rate, combined inform and success rate). This largely increase the difficulty to establish a suitable/fair standard for comparison. Also, some scenarios may require multiple metrics to determine overall performance (Samuel et al., 2024). For instance, we might need to assess the fluency, empathy, and safety of personalized LLMs. Consequently, we do not include a comprehensive evaluation comparison in the paper. Instead, we provide the solid taxonomy, content, and future directions that could serve as both a valuable introduction for newcomers and a comprehensive resource for current researchers in the field.
@@ Acknowledgements
We thank Yu-Ching Hsu and Jia-Yin Foo from National Taiwan University for their assistance and discussion. This work was financially supported by the National Science and Technology Council (NSTC) in Taiwan, under Grants 112-2223-E-002-012-MY5 and 111-$2$-E-002-013-MY3, and from the Featured Area Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education (113L900901/113L900902/113L900903).
@@ Appendix A Web
Liu et al. (2023a) Beauty Amazon Prompting Rating Prediction, Sequential Recommendation, Direct Recommendation, Explanation Generation and Review Summarization
@@
Chen et al. (2022a) News MIND Wu et al. (2020), Reddit Fine-tuning with weak labels Direct Recommendation
Table 2: An overview of existing research in recommendation. Following the classification of Liu et al. (2023a), we classify recommendation systems into five types: rating prediction, sequential recommendation, explanation Generation, and review generation, and direct recommendation.
@@
User Persona PersonaChat Zhang et al. (2018b) 11,907 164,356 1
@@
Table 4: A list of commonly used datasets for ToD modeling and user persona modeling. Among them, different versions of MultiWOZ Budzianowski et al. (2018); Ramadan et al. (2018); Eric et al. (2020); Zang et al. (2020) and PersonaChat Zhang et al. (2018b) are the most commonly used. Updated versions of MultiWOZ improve in several aspects: data quality, dialogue complexity, schema and ontology updates, and dataset sizes. PersonaChat contains various persona profiles, consisting of background, preferences, and personality traits. These profiles enable the modeling of coherent and contextual multi-turn diverse dialogue scenarios. For applications in user persona modeling, Tu et al. (2023) match individuals with persona-compatible virtual supporters and introduces the MBTI-S2Conv dataset, containing conversations between characters with distinct profiles. Lotfi et al. (2024) and Han et al. (2024) both propose synthetic datasets related to the Big Five personality.
@@
Table 6: An example of user persona modeling Section 3.5 from Persona-Chat dataset Zhang et al. (2018a).